OpenStudy (anonymous):

solve the equation algebraically 3(3x-1)^2=21

OpenStudy (anonymous):

\[3(3x-1)^{2}=21\]

OpenStudy (anonymous):

\[(3x-1)^{2}=7\]

OpenStudy (anonymous):

\[3x-1=\sqrt{7}\]

OpenStudy (anonymous):

\[3x=1+\sqrt{7}\]

OpenStudy (anonymous):

\[x=\frac{ 1+\sqrt{7} }{ 3 }\]

OpenStudy (anonymous):

don't forget the \(\pm\)

OpenStudy (anonymous):

\[x=\frac{ 1\pm\sqrt{7} }{ 3 }\]